An elevated level of sodium ions in the blood would be 0.155 mol L-1 and would be cause for concern. The average adult female has a total blood volume of 4.70 L. Calculate the total mass of sodium ions present when this concentration is found, assuming a blood volume of 4.70 L. 


1
Expert's answer
2022-03-27T10:28:03-0400

Number of moles per litre= 0.155

Total moles in an adult= 0.155×4.7= 0.7285

Mass= 23× 0.7285= 16.76g